# From Rankings to Insights: Evaluation Should Shift Focus from Leaderboard to Feedback

### v2

@@ C.1.3 Algorithm for TaxBuilder
11:     $\triangleright$ Child level: LLM returns parent node name
@@ C.1.4 Prompts for TaxBuilder
Prompt of LLM-as-Decision-Maker (Node Insertion) Your task is to add a new node to a tree-structured classification system. I will give you some [Current Level Nodes] and a [New Node]. You need to determine the next action and respond strictly in the specified format. Before giving the final answer, you must perform brief analysis.
# Task 1
Analyze whether the [New Node] shares identical meaning with any [Current Level Nodes] (including exact matches, case variations, abbreviations, or conceptually equivalent expressions). If identical, return <decision>EXIST</decision> and skip subsequent tasks.
Analyze whether the [New Node] represents a specific category or a classification criterion. If it’s a classification criterion, return <decision>EXIST</decision> and skip subsequent tasks.
# Task 2
Analyze whether the [New Node] doesn’t belong to any subordinate node of [Current Level Nodes] but should be at the same level. If yes, return <decision>ADD</decision> and skip subsequent tasks.
# Task 3
Identify which [Current Level Node] the [New Node] should belong to as a subordinate node. Return the parent node’s name wrapped in <decision></decision>.
## Example
Current Level Nodes: ["Fruits", "Vegetables", "Electronics"]
New Node: Phone
Output: <decision>Electronics</decision>
## New Node
{new_node}
## Current Level Nodes
{current_keys}
## Reference Classification System (for reference only)
{init_tree}
@@
Prompt of LLM-as-Decision-Maker (Node Refinement and Pruning) Your task is to optimize a given tree structure classification system. I will provide all the next-level nodes under a certain category (referred to as "current layer nodes") and the leaf nodes among them (referred to as "leaf nodes in the current layer"). You need to determine the next steps and return results in the specified format. Before giving the final answer, please provide a brief analysis. Note: Do not modify the names of the original categories, and do not omit any categories from the "current layer nodes" when returning results. Each task analysis should analyze all categories without omission.
If you see that the "current layer nodes" are all classification criteria rather than specific categories, this layer absolutely must not be modified.
### Task 1: Split
Examine all nodes in the "leaf nodes in the current layer" one by one, determine if there are any nodes that are not "single concepts" and split them. If the newly split category already exists, do not return that new category. A "single concept" refers to an independent concept, as opposed to a "combined concept" which is formed by combining multiple concepts. For example, "mythology novel" is a combined concept, consisting of "mythology" and "novel". Return format:
[(old_key1, [new_key1, new_key2]), (old_key2, [new_key1, new_key2])]
### Task 2: Parent-Child Relationship
Analyze all nodes in the "current layer nodes" one by one, determine if there exist parent-child relationships, or if multiple nodes can be merged into a parent node, and return the results. If there are no qualifying situations, do not return anything. Return format:
[(parent_key1, [old_child_key1, old_child_key2]), (parent_key2, [old_child_key1, old_child_key2])]
### Task 3: Delete
Examine all nodes in the "current layer nodes" one by one, identify meaningless categories (such as null values, blanks, etc.) or extremely rare categories in the real world, and delete them. Return format:
[old_key_to_delete1, old_key_to_delete2]
### Task 4: Merge
Analyze all nodes in the "current layer nodes" one by one, identify categories with consistent concepts and merge them. Return format:
[(new_key1, [old_key_to_merge1, old_key_to_merge2]), (new_key2, [old_key_to_merge1, old_key_to_merge2])]
### Task 5: Keep
Preserve the remaining nodes. Return format:
[old_key1, old_key2]
## Output Format
Provide a brief analysis for each task and give the results after the task is executed.
After completing all tasks, format the final result as a Python dictionary wrapped in <decision></decision> tags, in the following format, where all keys with old_ must have appeared in the "current layer nodes":
<decision> {
"split": [(old_key1, [new_key1, new_key2]), (old_key2, [new_key1, new_key2])],
"reparent": [(parent_key1, [old_child_key1, old_child_key2]), (parent_key2, [old_child_key1, old_child_key2])],
"delete": [old_key_to_delete1, old_key_to_delete2],
"merge": [(new_key1, [old_key_to_merge1, old_key_to_merge2]), (new_key2, [old_key_to_merge1, old_key_to_merge2])],
"keep": [old_key1, old_key2]
} </decision>
If there is no operation for a certain task, do not display that item in the final output.
## Current Layer Nodes
{current_keys}
## Leaf Nodes in the Current Layer
{leaf_keys}
@@ C.2.1 Seed Data
Prompt of Domain Annotation You are tasked with categorizing a given query into one of the following six categories:
1. **roleplay**
2. **coding**
3. **mathematics**
4. **reasoning**
5. **knowledge**
6. **writing**
7. **other**
For each query provided, determine the most appropriate category and output the result in lowcase enclosed within <domain> and </domain> tags.
**Example**:
Query: "How do I write a compelling essay introduction?"
Output: <domain>roleplay</domain>
**Now, analyze the following query**:
<|begin_of_query|>
{query}
<|end_of_query|>
@@
Prompt of Query Tags Annotation (writing) You are tasked with categorizing a given query into multiple tags based on the following hierarchical classification system.
### Classification System:
{taxonomy}
### Rules for Tagging:
1. **Hierarchy Rule**: If a query matches both a parent node and its child nodes, include both in the tags.
2. **Multiple Matches**: If a query matches multiple nodes at the same level, include all matching nodes.
3. **No Match**: If a query does not match any nodes under the second-level main categories, assign the tag "Other" to that category.
4. **Output Format**: The final output must be enclosed within ‘<tags>‘ and ‘</tags>‘ tags, and the tags should be provided as a JSON object where the keys are the basis for classification and the values are lists of matching tags.
### Example:
Query: "I need help brainstorming ideas for a fantasy story involving a hidden kingdom and magical creatures."
Output: <tags>{{"Creative Stages": ["Conceptualization Stage", "Brainstorming", "World Building", "Creative Exploration", "Content Conceptualization"], "Writing Domains": ["Literary Writing", "Fiction", "Fantasy", "Story Creation"], "Styles": ["Other"]}}</tags>
### Now, analyze the following query:
<|begin_of_query|>
{query}
<|end_of_query|>
@@
Prompt of Query Tags Annotation (roleplay) You are tasked with categorizing a given query into multiple tags based on the following hierarchical classification system.
### Classification System:
{taxonomy}
### Rules for Tagging:
1. **Hierarchy Rule**: If a query matches both a parent node and its child nodes, include both in the tags.
2. **Multiple Matches**: If a query matches multiple nodes at the same level, include all matching nodes.
3. **No Match**: If a query does not match any nodes under the second-level main categories, assign the tag "Other" to that category.
4. **Output Format**: The final output must be enclosed within ‘<tags>‘ and ‘</tags>‘ tags, and the tags should be provided as a JSON object where the keys are the basis for classification and the values are lists of matching tags.
### Example:
Query: "Write a script for a medieval fantasy story involving a knight’s adventure and a magical encounter, with a humorous twist."
Output: <tags>{{"Theme Types": ["Fantasy", "Medieval", "Magic", "Comedy", "Humorous Comedy"], "Task Types": ["Creative", "Scriptwriting"], "Style Types": ["Humorous Style", "Light Humor", "Narrative"]}}</tags>
### Now, analyze the following query:
<|begin_of_query|>
{query}
<|end_of_query|>
@@
Prompt of Query Tags Annotation (knowledge) You are tasked with categorizing a given query into multiple tags based on the following hierarchical classification system.
### Classification System:
{taxonomy}
### Rules for Tagging:
1. **Hierarchy Rule**: If a query matches both a parent node and its child nodes, include both in the tags.
2. **Multiple Matches**: If a query matches multiple nodes at the same level, include all matching nodes.
3. **No Match**: If a query does not match any nodes under the second-level main categories, assign the tag "Other" to that category.
4. **Output Format**: The final output must be enclosed within ‘<tags>‘ and ‘</tags>‘ tags, and the tags should be provided as a JSON object where the keys are the basis for classification and the values are lists of matching tags.
### Example:
Query: "How can I analyze data from a physics experiment on thermodynamics and present it effectively?"
Output: <tags>{{"Disciplinary Fields": ["Natural Sciences", "Physics", "Thermodynamics"], "Cognitive Levels": ["Applied Analysis"], "Task Types": ["Information Processing", "Data Analysis", "Content Production", "Content Generation"]}}</tags>
### Now, analyze the following query:
<|begin_of_query|>
{query}
<|end_of_query|>
@@
Prompt of Query Tags Annotation (coding) You are tasked with categorizing a given query into multiple tags based on the following hierarchical classification system.
### Classification System:
{taxonomy}
### Rules for Tagging:
1. **Hierarchy Rule**: If a query matches both a parent node and its child nodes, include both in the tags.
2. **Multiple Matches**: If a query matches multiple nodes at the same level, include all matching nodes.
3. **No Match**: If a query does not match any nodes under the second-level main categories, assign the tag "Other" to that category.
4. **Output Format**: The final output must be enclosed within ‘<tags>‘ and ‘</tags>‘ tags, and the tags should be provided as a JSON object where the keys are the basis for classification and the values are lists of matching tags.
### Example:
Query: "How can I optimize a Python script that processes large datasets and visualizes the results?"
Output: <tags>{{"Task Types": ["Data Processing", "Data Analysis", "Data Visualization", "Debugging and Optimization", "Optimization"], "Technical Domains": ["Data Science", "Data Analysis Methods", "Data Visualization Tools"], "Programming Languages": ["General-purpose Languages", "Python"]}}</tags>
### Now, analyze the following query:
<|begin_of_query|>
{query}
<|end_of_query|>
@@
Prompt of Query Tags Annotation (mathematics) You are tasked with categorizing a given query into multiple tags based on the following hierarchical classification system.
### Classification System:
{taxonomy}
### Rules for Tagging:
1. **Hierarchy Rule**: If a query matches both a parent node and its child nodes, include both in the tags.
2. **Multiple Matches**: If a query matches multiple nodes at the same level, include all matching nodes.
3. **No Match**: If a query does not match any nodes under the second-level main categories, assign the tag "Other" to that category.
4. **Output Format**: The final output must be enclosed within ‘<tags>‘ and ‘</tags>‘ tags, and the tags should be provided as a JSON object where the keys are the basis for classification and the values are lists of matching tags.
### Example:
Query: "How do I use calculus to model the growth of a population over time and graph the results?"
Output: <tags>{{"Mathematical Subfields": ["Analysis", "Calculus", "Applied Mathematics"], "Task Types": ["Problem Solving", "Modeling", "Visualization", "Function Graphing"]}}</tags>
### Now, analyze the following query:
<|begin_of_query|>
{query}
<|end_of_query|>
@@
Prompt of Query Tags Annotation (reasoning) You are tasked with categorizing a given query into multiple tags based on the following hierarchical classification system.
### Classification System:
{taxonomy}
### Rules for Tagging:
1. **Hierarchy Rule**: If a query matches both a parent node and its child nodes, include both in the tags.
2. **Multiple Matches**: If a query matches multiple nodes at the same level, include all matching nodes.
3. **No Match**: If a query does not match any nodes under the second-level main categories, assign the tag "Other" to that category.
4. **Output Format**: The final output must be enclosed within ‘<tags>‘ and ‘</tags>‘ tags, and the tags should be provided as a JSON object where the keys are the basis for classification and the values are lists of matching tags.
### Example:
Query: "How can I determine the cause of an event based on multiple contributing factors?"
Output: <tags>{{"Reasoning Methods": ["Multi-factor Attribution"], "Application Domains": ["Other"], "Thinking Modes": ["Analytical"], "Task Types": ["Reasoning"]}}</tags>
### Now, analyze the following query:
<|begin_of_query|>
{query}
<|end_of_query|>
@@
Prompt of Quality Annotation (writing) ## Task
Your task is to assess the quality of a given query based on the checklists outlined below. For each checklist, provide a clear explanation of your reasoning before assigning "Yes" or "No". If your answer is "Borderline", please answer "Yes" for this checklist.
## Checklists
Determine whether the provided question meets the following criteria. Return a Python array listing the numbers of all satisfied criteria:
1. **Clarity**: Is the question clear and well-defined?
2. **Completeness**: Does the question provide enough information for the LLM to answer the question?
3. **Complexity**: Does the question have enough depth and challenge beyond simple fact recall?
4. **Real-World Application**: Is this writing task something that would be proposed in the real world?
5. **Professionalism**: Does it require professional capabilities or professional knowledge?
6. **Originality:** Does the question encourage or require originality?
7. **User’s Requirements**: Does the user have clear, detailed, or unique requests that need to be considered in the response?
For example, if the question meets Clarity, Completeness, and Real-World Application, return ‘[1, 2, 4]‘.
## Final Output
For each question provided, return a Python dictionary in the following format:
“‘python
Final Labels: {{"question_quality": [1, 3, 4]}}
## The Query You Should Process
<|begin_of_query|>
{query}
<|end_of_query|>
@@
Prompt of Quality Annotation (roleplay) ## Task
Your task is to assess the quality of a given query based on the checklists outlined below. For each checklist, provide a clear explanation of your reasoning before assigning "Yes" or "No". If your answer is "Borderline", please answer "Yes" for this checklist.
## Checklists
Determine whether the provided question meets the following criteria. Return a Python array listing the numbers of all satisfied criteria:
1. **Clarity**: Is the question clear and well-defined?
2. **Completeness**: Does the question provide enough information for the LLM to answer the question?
3. **Complexity**: Does it involve in-depth understanding of any role-playing content, such as the psychology, characterization, and world-building of characters?
4. **Real-World Application**: Is this role-playing task something that would be proposed in the real world?
5. **Interactivity**: Does the question encourage meaningful interactions between characters, rather than single character?
6. **Engagement**: Does the task motivate active participation and emotional involvement from the audience or participants?
7. **Creativity:** Does it have creativity and novelty, or does solving it require creativity?
For example, if the question meets Clarity, Completeness, and Real-World Application, return ‘[1, 2, 4]‘.
## Final Output
For each question provided, return a Python dictionary in the following format:
“‘python
Final Labels: {{"question_quality": [1, 3, 4]}}
## The Query You Should Process
<|begin_of_query|>
{query}
<|end_of_query|>
@@
Prompt of Quality Annotation (knowledge) ## Task
Your task is to assess the quality of a given query based on the checklists outlined below. For each checklist, provide a clear explanation of your reasoning before assigning "Yes" or "No". If your answer is "Borderline", please answer "Yes" for this checklist.
## Checklists
Determine whether the provided question meets the following criteria. Return a Python array listing the numbers of all satisfied criteria:
1. **Clarity**: Is the question clear and well-defined?
2. **Completeness**: Does the question provide enough information for the LLM to answer the question?
3. **Complexity**: Does the question have enough depth and challenge beyond simple fact recall?
4. **Real-World Application**: Is the question something that would be encountered in real-world?
5. **Depth of Knowledge**: Does the question require deep expertise in the subject instead of just memory?
6. **Cross-Disciplinary**: Does the question involve cross-disciplinary aspects?
7. **Open-Endedness.**: Does the question encourage open-ended responses rather than simple “yes” or “no” answers, promoting deeper thinking?
For example, if the question meets Clarity, Completeness, and Real-World Application, return ‘[1, 2, 4]‘.
## Final Output
For each question provided, return a Python dictionary in the following format:
“‘python
Final Labels: {{"question_quality": [1, 3, 4]}}
## The Query You Should Process
<|begin_of_query|>
{query}
<|end_of_query|>
@@
Prompt of Quality Annotation (coding) ## Task
Your task is to assess the quality of a given query based on the checklists outlined below. For each checklist, provide a clear explanation of your reasoning before assigning "Yes" or "No". If your answer is "Borderline", please answer "No" for this checklist.
## Checklists
Determine whether the provided question meets the following criteria. Return a Python array listing the numbers of all satisfied criteria:
1. **Clarity**: Is the question clear and well-defined?
2. **Completeness**: Does the question provide enough information for the LLM to answer the question?
3. **Complexity**: Does it involve multiple components, layers, or nuance?
4. **Real-World Application**: Is the question something that would be encountered in real-world development?
5. **Problem-Solving**: Does it require active problem-solving beyond simple and superficial script or fact recall?
6. **Domain-Specific Expertise**: Does the question require in-depth knowledge of at least one specific area of programming?
7. **Specified Requirements**: Does it specify particular requirements, such as execution time, space constraints, specific programming language, tools, packages, etc.?
For example, if the question meets Clarity, Completeness, and Real-World Application, return ‘[1, 2, 4]‘.
## Final Output
For each question provided, return a Python dictionary in the following format:
“‘python
Final Labels: {{"question_quality": [1, 3, 4]}}
## The Query You Should Process
<|begin_of_query|>
{query}
<|end_of_query|>
@@
Prompt of Quality Annotation (mathematics) ## Task
Your task is to assess the query of a given question based on the rules outlined below. For each rule, provide a clear explanation of your reasoning before assigning a label.
## Rule
Determine whether the provided query meets the following criteria. Return a Python array listing the numbers of all satisfied criteria:
1. **Clarity**: Is the question clear and well-defined?
2. **Completeness**: Does the question provide enough information for the LLM to answer the question?
3. **Complexity**: Does it involve multiple steps, analysis, or reasoning instead of simple concept memorization and numerical calculation?
4. **Real-World Application**: Is the question something that would be encountered in real-world?
5. **Problem-Solving**: Does it test the ability to apply math in some scenarios?
6. **Rigorous Logic**: Does it involve content such as theorem derivation and formula understanding, which require rigorous logical abilities?
7. **Creativity:** Does it have creativity and novelty, or does solving it require creativity?
For example, if the question meets Clarity, Completeness, and Real-World Application, return ‘[1, 2, 4]‘.
## Final Output
For each question provided, return a Python dictionary in the following format:
“‘python
Final Labels: {{"question_quality": [1, 3, 4]}}
## The Query You Should Process
<|begin_of_query|>
{query}
<|end_of_query|>

### v3

@@ C.1.3 Algorithm for TaxBuilder
[…]
@@ C.1.4 Prompts for TaxBuilder
Prompt of LLM-as-Decision-Maker (Node Insertion) Your task is to add a new node to a tree-structured classification system. I will give you some [Current Level Nodes] and a [New Node]. You need to determine the next action and respond strictly in the specified format. Before giving the final answer, you must perform brief analysis. # Task 1 Analyze whether the [New Node] shares identical meaning with any [Current Level Nodes] (including exact matches, case variations, abbreviations, or conceptually equivalent expressions). If identical, return <decision>EXIST</decision> and skip subsequent tasks. Analyze whether the [New Node] represents a specific category or a classification criterion. If it’s a classification criterion, return <decision>EXIST</decision> and skip subsequent tasks. # Task 2 Analyze whether the [New Node] doesn’t belong to any subordinate node of [Current Level Nodes] but should be at the same level. If yes, return <decision>ADD</decision> and skip subsequent tasks. # Task 3 Identify which [Current Level Node] the [New Node] should belong to as a subordinate node. Return the parent node’s name wrapped in <decision></decision>. ## Example Current Level Nodes: ["Fruits", "Vegetables", "Electronics"] New Node: Phone Output: <decision>Electronics</decision> ## New Node {new_node} ## Current Level Nodes {current_keys} ## Reference Classification System (for reference only) {init_tree}
@@
Prompt of LLM-as-Decision-Maker (Node Refinement and Pruning) Your task is to optimize a given tree structure classification system. I will provide all the next-level nodes under a certain category (referred to as "current layer nodes") and the leaf nodes among them (referred to as "leaf nodes in the current layer"). You need to determine the next steps and return results in the specified format. Before giving the final answer, please provide a brief analysis. Note: Do not modify the names of the original categories, and do not omit any categories from the "current layer nodes" when returning results. Each task analysis should analyze all categories without omission. If you see that the "current layer nodes" are all classification criteria rather than specific categories, this layer absolutely must not be modified. ### Task 1: Split Examine all nodes in the "leaf nodes in the current layer" one by one, determine if there are any nodes that are not "single concepts" and split them. If the newly split category already exists, do not return that new category. A "single concept" refers to an independent concept, as opposed to a "combined concept" which is formed by combining multiple concepts. For example, "mythology novel" is a combined concept, consisting of "mythology" and "novel". Return format: [(old_key1, [new_key1, new_key2]), (old_key2, [new_key1, new_key2])] ### Task 2: Parent-Child Relationship Analyze all nodes in the "current layer nodes" one by one, determine if there exist parent-child relationships, or if multiple nodes can be merged into a parent node, and return the results. If there are no qualifying situations, do not return anything. Return format: [(parent_key1, [old_child_key1, old_child_key2]), (parent_key2, [old_child_key1, old_child_key2])] ### Task 3: Delete Examine all nodes in the "current layer nodes" one by one, identify meaningless categories (such as null values, blanks, etc.) or extremely rare categories in the real world, and delete them. Return format: [old_key_to_delete1, old_key_to_delete2] ### Task 4: Merge Analyze all nodes in the "current layer nodes" one by one, identify categories with consistent concepts and merge them. Return format: [(new_key1, [old_key_to_merge1, old_key_to_merge2]), (new_key2, [old_key_to_merge1, old_key_to_merge2])] ### Task 5: Keep Preserve the remaining nodes. Return format: [old_key1, old_key2] ## Output Format Provide a brief analysis for each task and give the results after the task is executed. After completing all tasks, format the final result as a Python dictionary wrapped in <decision></decision> tags, in the following format, where all keys with old_ must have appeared in the "current layer nodes": <decision> { "split": [(old_key1, [new_key1, new_key2]), (old_key2, [new_key1, new_key2])], "reparent": [(parent_key1, [old_child_key1, old_child_key2]), (parent_key2, [old_child_key1, old_child_key2])], "delete": [old_key_to_delete1, old_key_to_delete2], "merge": [(new_key1, [old_key_to_merge1, old_key_to_merge2]), (new_key2, [old_key_to_merge1, old_key_to_merge2])], "keep": [old_key1, old_key2] } </decision> If there is no operation for a certain task, do not display that item in the final output. ## Current Layer Nodes {current_keys} ## Leaf Nodes in the Current Layer {leaf_keys}
@@ C.2.1 Seed Data
Prompt of Domain Annotation You are tasked with categorizing a given query into one of the following six categories: 1. **roleplay** 2. **coding** 3. **mathematics** 4. **reasoning** 5. **knowledge** 6. **writing** 7. **other** For each query provided, determine the most appropriate category and output the result in lowcase enclosed within <domain> and </domain> tags. **Example**: Query: "How do I write a compelling essay introduction?" Output: <domain>roleplay</domain> **Now, analyze the following query**: <|begin_of_query|> {query} <|end_of_query|>
@@
Prompt of Query Tags Annotation (writing) You are tasked with categorizing a given query into multiple tags based on the following hierarchical classification system. ### Classification System: {taxonomy} ### Rules for Tagging: 1. **Hierarchy Rule**: If a query matches both a parent node and its child nodes, include both in the tags. 2. **Multiple Matches**: If a query matches multiple nodes at the same level, include all matching nodes. 3. **No Match**: If a query does not match any nodes under the second-level main categories, assign the tag "Other" to that category. 4. **Output Format**: The final output must be enclosed within ‘<tags>‘ and ‘</tags>‘ tags, and the tags should be provided as a JSON object where the keys are the basis for classification and the values are lists of matching tags. ### Example: Query: "I need help brainstorming ideas for a fantasy story involving a hidden kingdom and magical creatures." Output: <tags>{{"Creative Stages": ["Conceptualization Stage", "Brainstorming", "World Building", "Creative Exploration", "Content Conceptualization"], "Writing Domains": ["Literary Writing", "Fiction", "Fantasy", "Story Creation"], "Styles": ["Other"]}}</tags> ### Now, analyze the following query: <|begin_of_query|> {query} <|end_of_query|>
@@
Prompt of Query Tags Annotation (roleplay) You are tasked with categorizing a given query into multiple tags based on the following hierarchical classification system. ### Classification System: {taxonomy} ### Rules for Tagging: 1. **Hierarchy Rule**: If a query matches both a parent node and its child nodes, include both in the tags. 2. **Multiple Matches**: If a query matches multiple nodes at the same level, include all matching nodes. 3. **No Match**: If a query does not match any nodes under the second-level main categories, assign the tag "Other" to that category. 4. **Output Format**: The final output must be enclosed within ‘<tags>‘ and ‘</tags>‘ tags, and the tags should be provided as a JSON object where the keys are the basis for classification and the values are lists of matching tags. ### Example: Query: "Write a script for a medieval fantasy story involving a knight’s adventure and a magical encounter, with a humorous twist." Output: <tags>{{"Theme Types": ["Fantasy", "Medieval", "Magic", "Comedy", "Humorous Comedy"], "Task Types": ["Creative", "Scriptwriting"], "Style Types": ["Humorous Style", "Light Humor", "Narrative"]}}</tags> ### Now, analyze the following query: <|begin_of_query|> {query} <|end_of_query|>
@@
Prompt of Query Tags Annotation (knowledge) You are tasked with categorizing a given query into multiple tags based on the following hierarchical classification system. ### Classification System: {taxonomy} ### Rules for Tagging: 1. **Hierarchy Rule**: If a query matches both a parent node and its child nodes, include both in the tags. 2. **Multiple Matches**: If a query matches multiple nodes at the same level, include all matching nodes. 3. **No Match**: If a query does not match any nodes under the second-level main categories, assign the tag "Other" to that category. 4. **Output Format**: The final output must be enclosed within ‘<tags>‘ and ‘</tags>‘ tags, and the tags should be provided as a JSON object where the keys are the basis for classification and the values are lists of matching tags. ### Example: Query: "How can I analyze data from a physics experiment on thermodynamics and present it effectively?" Output: <tags>{{"Disciplinary Fields": ["Natural Sciences", "Physics", "Thermodynamics"], "Cognitive Levels": ["Applied Analysis"], "Task Types": ["Information Processing", "Data Analysis", "Content Production", "Content Generation"]}}</tags> ### Now, analyze the following query: <|begin_of_query|> {query} <|end_of_query|>
@@
Prompt of Query Tags Annotation (coding) You are tasked with categorizing a given query into multiple tags based on the following hierarchical classification system. ### Classification System: {taxonomy} ### Rules for Tagging: 1. **Hierarchy Rule**: If a query matches both a parent node and its child nodes, include both in the tags. 2. **Multiple Matches**: If a query matches multiple nodes at the same level, include all matching nodes. 3. **No Match**: If a query does not match any nodes under the second-level main categories, assign the tag "Other" to that category. 4. **Output Format**: The final output must be enclosed within ‘<tags>‘ and ‘</tags>‘ tags, and the tags should be provided as a JSON object where the keys are the basis for classification and the values are lists of matching tags. ### Example: Query: "How can I optimize a Python script that processes large datasets and visualizes the results?" Output: <tags>{{"Task Types": ["Data Processing", "Data Analysis", "Data Visualization", "Debugging and Optimization", "Optimization"], "Technical Domains": ["Data Science", "Data Analysis Methods", "Data Visualization Tools"], "Programming Languages": ["General-purpose Languages", "Python"]}}</tags> ### Now, analyze the following query: <|begin_of_query|> {query} <|end_of_query|>
@@
Prompt of Query Tags Annotation (mathematics) You are tasked with categorizing a given query into multiple tags based on the following hierarchical classification system. ### Classification System: {taxonomy} ### Rules for Tagging: 1. **Hierarchy Rule**: If a query matches both a parent node and its child nodes, include both in the tags. 2. **Multiple Matches**: If a query matches multiple nodes at the same level, include all matching nodes. 3. **No Match**: If a query does not match any nodes under the second-level main categories, assign the tag "Other" to that category. 4. **Output Format**: The final output must be enclosed within ‘<tags>‘ and ‘</tags>‘ tags, and the tags should be provided as a JSON object where the keys are the basis for classification and the values are lists of matching tags. ### Example: Query: "How do I use calculus to model the growth of a population over time and graph the results?" Output: <tags>{{"Mathematical Subfields": ["Analysis", "Calculus", "Applied Mathematics"], "Task Types": ["Problem Solving", "Modeling", "Visualization", "Function Graphing"]}}</tags> ### Now, analyze the following query: <|begin_of_query|> {query} <|end_of_query|>
@@
Prompt of Query Tags Annotation (reasoning) You are tasked with categorizing a given query into multiple tags based on the following hierarchical classification system. ### Classification System: {taxonomy} ### Rules for Tagging: 1. **Hierarchy Rule**: If a query matches both a parent node and its child nodes, include both in the tags. 2. **Multiple Matches**: If a query matches multiple nodes at the same level, include all matching nodes. 3. **No Match**: If a query does not match any nodes under the second-level main categories, assign the tag "Other" to that category. 4. **Output Format**: The final output must be enclosed within ‘<tags>‘ and ‘</tags>‘ tags, and the tags should be provided as a JSON object where the keys are the basis for classification and the values are lists of matching tags. ### Example: Query: "How can I determine the cause of an event based on multiple contributing factors?" Output: <tags>{{"Reasoning Methods": ["Multi-factor Attribution"], "Application Domains": ["Other"], "Thinking Modes": ["Analytical"], "Task Types": ["Reasoning"]}}</tags> ### Now, analyze the following query: <|begin_of_query|> {query} <|end_of_query|>
@@
Prompt of Quality Annotation (writing) ## Task Your task is to assess the quality of a given query based on the checklists outlined below. For each checklist, provide a clear explanation of your reasoning before assigning "Yes" or "No". If your answer is "Borderline", please answer "Yes" for this checklist. ## Checklists Determine whether the provided question meets the following criteria. Return a Python array listing the numbers of all satisfied criteria: 1. **Clarity**: Is the question clear and well-defined? 2. **Completeness**: Does the question provide enough information for the LLM to answer the question? 3. **Complexity**: Does the question have enough depth and challenge beyond simple fact recall? 4. **Real-World Application**: Is this writing task something that would be proposed in the real world? 5. **Professionalism**: Does it require professional capabilities or professional knowledge? 6. **Originality:** Does the question encourage or require originality? 7. **User’s Requirements**: Does the user have clear, detailed, or unique requests that need to be considered in the response? For example, if the question meets Clarity, Completeness, and Real-World Application, return ‘[1, 2, 4]‘. ## Final Output For each question provided, return a Python dictionary in the following format: “‘python Final Labels: {{"question_quality": [1, 3, 4]}} “‘ ## The Query You Should Process <|begin_of_query|> {query} <|end_of_query|>
@@
Prompt of Quality Annotation (roleplay) ## Task Your task is to assess the quality of a given query based on the checklists outlined below. For each checklist, provide a clear explanation of your reasoning before assigning "Yes" or "No". If your answer is "Borderline", please answer "Yes" for this checklist. ## Checklists Determine whether the provided question meets the following criteria. Return a Python array listing the numbers of all satisfied criteria: 1. **Clarity**: Is the question clear and well-defined? 2. **Completeness**: Does the question provide enough information for the LLM to answer the question? 3. **Complexity**: Does it involve in-depth understanding of any role-playing content, such as the psychology, characterization, and world-building of characters? 4. **Real-World Application**: Is this role-playing task something that would be proposed in the real world? 5. **Interactivity**: Does the question encourage meaningful interactions between characters, rather than single character? 6. **Engagement**: Does the task motivate active participation and emotional involvement from the audience or participants? 7. **Creativity:** Does it have creativity and novelty, or does solving it require creativity? For example, if the question meets Clarity, Completeness, and Real-World Application, return ‘[1, 2, 4]‘. ## Final Output For each question provided, return a Python dictionary in the following format: “‘python Final Labels: {{"question_quality": [1, 3, 4]}} “‘ ## The Query You Should Process <|begin_of_query|> {query} <|end_of_query|>
@@
Prompt of Quality Annotation (knowledge) ## Task Your task is to assess the quality of a given query based on the checklists outlined below. For each checklist, provide a clear explanation of your reasoning before assigning "Yes" or "No". If your answer is "Borderline", please answer "Yes" for this checklist. ## Checklists Determine whether the provided question meets the following criteria. Return a Python array listing the numbers of all satisfied criteria: 1. **Clarity**: Is the question clear and well-defined? 2. **Completeness**: Does the question provide enough information for the LLM to answer the question? 3. **Complexity**: Does the question have enough depth and challenge beyond simple fact recall? 4. **Real-World Application**: Is the question something that would be encountered in real-world? 5. **Depth of Knowledge**: Does the question require deep expertise in the subject instead of just memory? 6. **Cross-Disciplinary**: Does the question involve cross-disciplinary aspects? 7. **Open-Endedness.**: Does the question encourage open-ended responses rather than simple “yes” or “no” answers, promoting deeper thinking? For example, if the question meets Clarity, Completeness, and Real-World Application, return ‘[1, 2, 4]‘. ## Final Output For each question provided, return a Python dictionary in the following format: “‘python Final Labels: {{"question_quality": [1, 3, 4]}} “‘ ## The Query You Should Process <|begin_of_query|> {query} <|end_of_query|>
@@
Prompt of Quality Annotation (coding) ## Task Your task is to assess the quality of a given query based on the checklists outlined below. For each checklist, provide a clear explanation of your reasoning before assigning "Yes" or "No". If your answer is "Borderline", please answer "No" for this checklist. ## Checklists Determine whether the provided question meets the following criteria. Return a Python array listing the numbers of all satisfied criteria: 1. **Clarity**: Is the question clear and well-defined? 2. **Completeness**: Does the question provide enough information for the LLM to answer the question? 3. **Complexity**: Does it involve multiple components, layers, or nuance? 4. **Real-World Application**: Is the question something that would be encountered in real-world development? 5. **Problem-Solving**: Does it require active problem-solving beyond simple and superficial script or fact recall? 6. **Domain-Specific Expertise**: Does the question require in-depth knowledge of at least one specific area of programming? 7. **Specified Requirements**: Does it specify particular requirements, such as execution time, space constraints, specific programming language, tools, packages, etc.? For example, if the question meets Clarity, Completeness, and Real-World Application, return ‘[1, 2, 4]‘. ## Final Output For each question provided, return a Python dictionary in the following format: “‘python Final Labels: {{"question_quality": [1, 3, 4]}} “‘ ## The Query You Should Process <|begin_of_query|> {query} <|end_of_query|>
@@
Prompt of Quality Annotation (mathematics) ## Task Your task is to assess the query of a given question based on the rules outlined below. For each rule, provide a clear explanation of your reasoning before assigning a label. ## Rule Determine whether the provided query meets the following criteria. Return a Python array listing the numbers of all satisfied criteria: 1. **Clarity**: Is the question clear and well-defined? 2. **Completeness**: Does the question provide enough information for the LLM to answer the question? 3. **Complexity**: Does it involve multiple steps, analysis, or reasoning instead of simple concept memorization and numerical calculation? 4. **Real-World Application**: Is the question something that would be encountered in real-world? 5. **Problem-Solving**: Does it test the ability to apply math in some scenarios? 6. **Rigorous Logic**: Does it involve content such as theorem derivation and formula understanding, which require rigorous logical abilities? 7. **Creativity:** Does it have creativity and novelty, or does solving it require creativity? For example, if the question meets Clarity, Completeness, and Real-World Application, return ‘[1, 2, 4]‘. ## Final Output For each question provided, return a Python dictionary in the following format: “‘python Final Labels: {{"question_quality": [1, 3, 4]}} “‘ ## The Query You Should Process <|begin_of_query|> {query} <|end_of_query|>